OpenStudy (anonymous):

A satellite of Jupiter has an orbital period of 1.77 days and an orbital of 4.22x10^5 km. Determine the mass of Jupiter. The answer is 1.90x10^27 kg but I do not know how to solve it.

OpenStudy (anonymous):

should the orbit be considered elliptical or circle?

OpenStudy (anonymous):

this is what i think could be done..if the orbital is circular its length is the circumference 2*pi*r and time period=2*pi*r/orbital velocity ....from this orbital velocity could be calculated ...we know orbital velocity is v=sqrt(GM/r) then calculate mass

OpenStudy (shane_b):

\[m = \frac{4pi^2r^3}{GT^2}\]\[m = \frac{4pi^2(4.22x10^8m)^3}{[6.674x10^{-11}N(m/kg)^2][1.77d*\frac{86400s}{d}]^2}=1.90x10^{27}kg\]Where: m = mass of the larger object in kg G = gravitational constant T = orbital period of satellite (in seconds) r = orbital distance of satellite from planet in meters

OpenStudy (anonymous):

Sorry I typed the problem wrong. It was supposed to say and an orbital RADIUS of 4.22x10^5 km. Therefore, the orbit is a circle. Sorry about that

OpenStudy (shane_b):

I assumed it was orbital radius before I did it :)

OpenStudy (anonymous):

Thank you :) When you plugged in for r did you mean to write 4.22x10^5 because you wrote 4.22x10^8?

OpenStudy (shane_b):

No...since the info you were given was in km you have to convert it to meters. I just did it on-the-fly there.

OpenStudy (shane_b):

It must be converted to meters for the formula...so that's why I basically multiplied it times 1000.

OpenStudy (shane_b):

Clear now or should I explain it better?

OpenStudy (anonymous):

oh ok i get it now! Thanks so much :)

OpenStudy (shane_b):

No problem, good luck !

OpenStudy (anonymous):

so is my answer correct @Shane_B and @Cyn_A